OpenStudy (anonymous):

what is d/dx of sqr root (1−x2−y2)?

OpenStudy (anonymous):

Rupesh get this it is from microsoft and will solve everything for you http://www.microsoft.com/download/en/details.aspx?id=15702

OpenStudy (anonymous):

\[\frac{-x}{\sqrt{-x^2-y^2+1}}\]

OpenStudy (anonymous):

= -(-2x)/ 2sqrt(1−x2−y2) = x/sqrt(1−x2−y2)

OpenStudy (anonymous):

Better: buy a ti-89 and it will even do linear algebra for you!

OpenStudy (anonymous):

agadgd that program will do linear too

OpenStudy (amistre64):

\[d/dx \sqrt{1−x^2−y^2}\] \[\frac{d/dx(1−x^2−y^2)}{\sqrt{1−x^2−y^2}}\] \[\frac{−2x−2y\ y'}{\sqrt{1−x^2−y^2}}\]

OpenStudy (amistre64):

forgot the 2 on the bottom :)

OpenStudy (amistre64):

\[\frac{−2x−2y\ y'}{2\sqrt{1−x^2−y^2}}\] \[\frac{2(−x−y\ y')}{2\sqrt{1−x^2−y^2}}\] \[\frac{−x−y\ y'}{\sqrt{1−x^2−y^2}}\] and solve for y'

OpenStudy (anonymous):

But y′ is zero, since we are differentiating with respect to x. Isnt it?

OpenStudy (amistre64):

no; y' is unknown

OpenStudy (amistre64):

but to clarify; is this partial derivative or implicit?

OpenStudy (anonymous):

Implicit

OpenStudy (amistre64):

then y' pops out; dy/dx doesnt = 0 all on its own

OpenStudy (amistre64):

its simply a rate of change in y with respect to x

OpenStudy (anonymous):

ok. I understand

OpenStudy (anonymous):

But my doubt started when I was trying to find the outward normal vector to a sphere. Can you help me with it?

OpenStudy (amistre64):

\[\frac{d}{dx} \sqrt{1−x^2−[y(x)]^2}\] \[\frac{{d}{dx}(1−x^2−[y(x)]^2)}{2\sqrt{1−x^2−[y(x)]^2}}\] \[\frac{\frac{d}{dx}1−\frac{d}{dx}x^2−\frac{d}{dx}[y(x)]^2}{2\sqrt{1−x^2−[y(x)]^2}}\] \[\frac{−2x−2[y(x)]*\frac{dy}{dx}}{2\sqrt{1−x^2−[y(x)]^2}}\] \[\frac{2(−x−[y(x)]*\frac{dy}{dx})}{2\sqrt{1−x^2−[y(x)]^2}}\] \[\frac{−x−[y(x)]*\frac{dy}{dx}}{\sqrt{1−x^2−[y(x)]^2}}\] \[\frac{−x−y*\frac{dy}{dx}}{\sqrt{1−x^2−y^2}}\]

OpenStudy (amistre64):

outward normal vector to a sphere is the partial of df/dx and df/dy

OpenStudy (anonymous):

agreed. But in my previous question on that , somebody replied that to be the grad of that function.

OpenStudy (anonymous):

But grad of that function and the the partial of df/dx and df/dy are turning out to be different. So I am confused?

OpenStudy (amistre64):

the gradient is the "equation" for the normal; same thing

OpenStudy (anonymous):

Sorry for taking too much of your time. In the question it was to find the direction of the outward normal to x2 + y2 + z2=1 at (1,1,1)? Kindly refer my previous question. Somebody replied me to take the grad of the function, which got to be 2 x r (where r is the radius vector) But when we take in parametric form by taking partial differentiation, are we getting the same answer. Confused. kindly help..

OpenStudy (amistre64):

\[f(x,y)= \sqrt{1−x^2−y^2}\] \[df/dx= \frac{-x}{\sqrt{1−x^2−y^2}}\] \[df/dy= \frac{-y}{\sqrt{1−x^2−y^2}}\] partials are done by treating the non-variable a constant the normal is: <df/fx, df/dy> for any given (x,y)

OpenStudy (amistre64):

internet is goofing up on my end

OpenStudy (amistre64):

if we take the partials of the f(x,y,z) the are simliar and the same concepts are involved

OpenStudy (amistre64):

the gradient provides general equations that are used to solve for the normal at any given point; so if we know the gradient, and a given point; we can know the normal at the point ...

OpenStudy (anonymous):

Oh. Ok.

OpenStudy (amistre64):



OpenStudy (amistre64):

"In the question it was to find the direction of the outward normal to x2 + y2 + z2=1 at (1,1,1)" \[f(x,y,z) = x^2 + y^2 + z^2 -1\] \[df/dx = 2x\] \[df/dx = 2y\] \[df/dx = 2z\] normal at (1,1,1) when the gradient is: 2<x,y,z> = 2<1,1,1>

OpenStudy (anonymous):

Okay. Much better understanding now.

OpenStudy (amistre64):

it just so happens that the vector from the origin (0,0,0) to the point (1,1,1) is: <1,1,1> and in sphere, the normal is in a straight line with the vector from the origin :) if that makes sense

OpenStudy (anonymous):

Okay. Much better understanding now.

OpenStudy (anonymous):

Thanks very much. Highly appreciated.

OpenStudy (amistre64):

youre welcome :)